OpenStudy (anonymous):

Simplify the functions and determine whether the functions are power functions. If a function is a power function, write it in a form y=ax^b a. y=x/3 b. y=1/3x

Directrix (directrix):

Grouping symbols for these equations would be helpful. Here's how I am reading the posted equations: a. y=x/3 --> y = (1/3) x b. y=1/3x --> y = (1/3) x or maybe y = 1/(3x) Please clarify. Thanks. @malibugranprix2000

OpenStudy (cwrw238):

if b is 1/(3x) this is a power function because it can be written in the form (1/3) x^-1

OpenStudy (anonymous):

@Directrix